1/4

Дмитрий Николаевич Садовников, Дмитрий Николаевич Садовников, Дмитрий Николаевич Садовников, Дмитрий Николаевич Садовников,

Volga, Volga (Stenka Rasin)

Stenka Rasin Sten’ka Razin Из-за острова на стрежень (Стенька Разин)

tradukita de N. N. 20  tradukita de N. N. 21  tradukita de Tr. Jacob Robbins

Bag om øen, på den brede underskønne Volgaflod Stenjka Rasins stolte både ud på nye togter stod.

In den Wellen hinter Inseln Ziehen Kähne malerisch, - |: Fangen leis an aufzuwachen, Bang ist jedes Angesicht. :| 1. From beyond the wooded island To the river wide and free Proudly sailed the arrow-breasted Ships of Cossack yeomanry.

1. Из-за острова на стрежень, На простор речной волны, Выплывают расписные, Острогрудые челны.

På den første høvding Stenjka sidder selv i bådens stavn, han på kamp og strid ej tænker, har prinsessen i sin favn.

Stenka Rasin vorn als erster, Selig in der Trunkenheit, |: Hält im Arme die Prinzessin, Die er eben erst befreit. :| 2. On the first is Stenka Razin With a princess at his side, |: Drunken, holds a marriage revel With his beautiful young bride. :| 2. На переднем Стенька Разин, Обнявшись, сидит с княжной, Свадьбу новую справляет, Сам веселый и хмельной.

Men hans mænd er Donkosakker, de er vrede, for de ved, han har glemt dem for en kvinde, for en nat i kærlighed.

.................................................. .................................................. .................................................. ..................................................

3. But behind them rose a whisper, “He has left his sword to woo; |: One short night, and Stenka Razin Has become a woman too!” :| 3. А она, потупив очи, Ни жива и ни мертва, Молча слушает хмельные Атамановы слова.

Stenjka Rasin hører latter, blodet koger i hans bryst Over Volgas brede strømme højt nu lyder Stenjkas røst.

.................................................. .................................................. .................................................. ..................................................

4. Stenka Razin hears the jeering Of his discontented band, |: And the lovely Persian princess He has circled with his hand. :| 4. Позади их слышен ропот:

- Нас на бабу променял! Только ночь с ней провозилс Сам наутро бабой стал .

…  …  …  …

2/4

....................................... ....................................... ....................................... .......................................

.................................................. .................................................. .................................................. ..................................................

5. His black brows have come together As the waves of anger rise, |: And the blood comes rushing swiftly To his piercing, jet-black eyes. :| 5. Этот ропот и насмешки Слышит грозный атаман, И могучею рукою Обнял персиянки стан.

....................................... ....................................... ....................................... .......................................

.................................................. .................................................. .................................................. ..................................................

6. “I will give you all you ask for, Life and heart, and head and hand,” |: Echo rolls the pealing thunder Of his voice across the land. :| 6. Брови черные сошлися, Надвигается гроза. Буйной кровью налилис Атамановы глаза.

....................................... ....................................... ....................................... .......................................

Plötzlich tönt ein dumpf Gemur-

re: Er verrät uns um ein Weib, |: All der Seinen Glück vergißt er Um geringen Zeitvertreib. :| 7. “Volga, Volga, mother Volga, Deep and wide beneath the sun, |: You have never seen a present From the Cossack of the Don. :| 7. ”Ничего не пожалею, Буйну голову отдам!“ — Раздается голос властный По окрестным берегам.

....................................... ....................................... ....................................... .......................................

Wolga, Wolga, Mutter Wolga, Ruhig fließest du dahin, |: Ahnst nicht, was ein Donkosake Fürchterliches hat im Sinn. :| 8. And that peace might rule as always All my free-born men and brave, |: Volga, Volga, mother Volga, Volga, make this girl a grave.” :| 8. “Волга, Волга, мать родная, Волга, русская река, Не видала ты подарка От донского казака!

”For at ingen splid skal være mellem frie mænd,- tag så Moder Volga, denne kvinde!“ og hans arm om hende lå.

.................................................. .................................................. .................................................. ..................................................

9. With a sudden, mighty movement, Razin lifts the beauty high, |: And he casts her where the waters Of the Volga move and sigh. :| 9. Чтобы не было раздора Между вольными людьми, Волга, Волга, мать родная, На, красавицу возьми!”

Og han ta’r med stærke arme da sin brud og kaster ned /: i de dybe, mørke bølger, sine nætters salighed. :/ Vorn als erster Stenka Rasin Hebt das Weib in wilder Wut, - |: Wolga, Wolga, nimm ein Opfer! Und er wirft sie in die Flut. :| 10. Now a silence like the grave sinks To all those who stand to see, |: And the battle-hardened Cossacks Sink to weep on bended knee. :| 10. Мощным взмахом поднимает Он красавицу княжну И за борт ее бросает В набежавшую волну.

…  …  …  …

3/4

“Men hvad fanden, hvorfor

hænger I med næbbet? Bliv kun ved! Dans og syng de røversange, syng for hendes sjælefred!”

Und er sieht sie untergehen, Hört noch ihren Jammerschrei: |: Stirb als Opfer meiner Treue! Stenka Rasin, er ist frei. :| 11. “Dance, you fool, and men, make mer-

ry! What has got into your eyes? |: Let us thunder out a chanty Of a place where beauty lies.” :| 11. “Что ж вы, братцы, приуныли? Эй, ты, Филька, черт, пляши! Грянем песню удалую На помин ее души!..”

Frem fra øen - ud på dybet stævner Stenjka Rasin ud /: med den stærke, stolte både, sejler fra sin døde brud. :/ Und die Kähne ziehen weiter, Und die Kähne ziehen fort, |: Und die Wolga fließet weiter Über diesen Schreckensort.

12. From beyond the wooded island

To the river wide and free,

|: Proudly sail the arrow-breasted Ships of Cossack yeomanry. :| 12. Из-за острова на стрежень, На простор речной волны, Выплывают расписные Острогрудые челны.

…  …  …  …

4/4

Traduko de la Rusa poemo “Из-за острова на стрежень (Стенька Разин)” de Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883) en la Danan de N. N. 20.

Arg-478-1037 (2010-03-08 12:16:20)

Aperinte en 1947 en la kant-libro “DLEA-Kantaro” eldonita de Dana Laborista Esperanto-Asocio (DLEA), n-ro 100 kaj en la kantlibro “Store Esperanto-Sangbog”, SEFO 1939, n-ro 31. Vidu la retejon esperantonia. dk sub http: // www. esperantonia. dk/ volga. htm .

Traduko de la Rusa poemo “Из-за острова на стрежень (Стенька Разин)” de Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883) en la Germanan de N. N. 21.

Arg-478-1038 (2010-03-09 15:31:43)

Mi, Manfred Retzlaff, trovis tiun ĉi germanigon de la kanto en http: // www. musicanet. org/ robokopp/ Lieder/ stenkara. html . Latinskriba aliskribo kaj laŭvorta germanigo troveblas en http: // www. kaikracht. de/ balalaika/ songs/ izza_bal. htm . Vidu ankaŭ la vikipedian retejon http: // de. wikipedia. org/ wiki/ Stenka_Rasin .

Traduko de la Rusa poemo “Из-за острова на стрежень (Стенька Разин)” de Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883) en la Anglan de Tr. Jacob Robbins en 1921.

Arg-478-1039 (2010-03-09 17:06:16)

Tiu ĉi angligo troviĝas en http: // ingeb. org/ songs/ stenkara. html .

Verkinto de tiu ĉi Rusa poemo estas Дмитрий Николаевич Садовников (Dmitrij Nikolajeviĉ Sadownikow, ∗1847 – †1883).

Arg-478-927 (2010-03-07 18:59:55)

Pri la aŭtoro vidu ankaŭ la informojn sub http: // ru. wikipedia. org , la ruslingva parto de Wikipedia. Informoj pri la kanto, la melodio kaj diversaj versioj de la kanto estas troveblaj en la retejo http: // a-pesni. golosa. info/ popular20/ izzaostrova. htm .